OpenStudy (anonymous):

Which function below is the inverse of f(x) = The quantity of three x plus five, over seven.?

OpenStudy (anonymous):

a)−1(x) = The quantity of seven x minus five, over three b) f−1(x) = The quantity of seven x plus five, over three C)−f−1(x) = Seven over the quantity of three x minus five d) f−1(x) = Seven over the quantity of five minus three x